### v1

@@ Abstract
A framework for Chebyshev spectral collocation methods for the numerical solution of functional and delay differential equations (FDEs and DDEs) is described. The framework combines interpolation via the barycentric resampling matrix with a multidomain approach used to resolve isolated discontinuities propagated by non-smooth initial data. Geometric convergence is demonstrated for several examples of linear and nonlinear FDEs and DDEs with various delay types, including discrete, proportional, continuous, and state-dependent delay. The framework is a natural extension of standard spectral collocation methods based on polynomial interpolants and can be readily incorporated into existing spectral discretisations, such as in Chebfun/Chebop, allowing the automated and efficient solution of a wide class of nonlinear functional and delay differential equations.
@@ 1 Introduction
Delay differential equations (DDEs) arise in a variety of application areas, notably population dynamics (Gopalsamy, 1992; Kuang, 1993), epidemiology (Brauer and Castillo-Chávez, 2001; Rihan, 2021), control theory (Kyrychko and Hogan, 2010; Michiels and Niculescu, 2007), and more (Erneux, 2009). Below we review some basic concepts and refer the reader to any of the many excellent introductory texts on DDEs and FDEs for details (Driver, 2012; Hale, 1971; Kolmanovskii and Myshkis, 1999; Michiels and Niculescu, 2007)).
The most elementary DDEs are first-order initial value problems of the form
@@
Like ordinary differential equations (ODEs), most DDEs and FDEs arising in practical settings do not have closed-form solutions, so some kind of approximation is required. Standard numerical methods for the approximating DDE solutions typically involve Runge–Kutta-like solvers equipped with a suitable interpolation to evaluate the delay terms (the method of steps), or rewriting the DDE as a system of ODEs with additional variables to represent the delayed terms. For an overview of such methods, we refer the reader to (Bellen and Zennaro, 2013), (Engelborghs et al., 2002), or (Shampine and Thompson, 2009), and references therein. A major difficulty arising in DDEs and FDEs not present in ODEs is the propagation of discontinuities arising from non-smooth initial data. Locating and resolving these discontinuities is necessary in order to compute an accurate and efficient approximation, even for low order methods (Baker et al., 1995).
@@
For smooth ODEs, spectral collocation methods (also known as pseudospectral methods) typically exhibit geometric convergence as the number of degrees of freedom is increased, making them particularly efficient when a high degree of accuracy is required (Boyd, 2001; Fornberg, 1996; Trefethen, 2000). Moreover, spectral collocation methods extend naturally to boundary value problems and differential-eigenvalue problems. Given the power of spectral methods in solving ODEs, it is natural to apply them to DDEs and FDEs, and there are numerous examples of this in the literature. We mention a few key papers here, but more detailed overviews of the development of spectral methods in these areas are given by Meng et al. (2019) and Jafari et al. (2021).
@@
One of the first spectral methods for DDEs was introduced by Fox et al. (1971), where a Chebyshev-tau approach was used to solve the linear pantograph equation. Later, a Legendre-tau approach for DDEs with discrete delay was presented by Ito et al. (1991). More recently there have been developments in using spectral methods to solve continuous delay DDEs (Brunner and Hu, 2005, 2007), nonlinear DDEs (Ali et al., 2009), neutral and state-dependent DDEs (Sedaghat et al., 2012), DDE differential eigenvalue problems (Liu and Ortiz, 1989), and periodic DDEs (Engelborghs et al., 2001; Butcher and Bobrenkov, 2011). In addition, spectral collocation methods have proven useful for computing characteristic roots of DDEs (Breda et al., 2005; Wu and Michiels, 2012). The work most closely related to that which we present here is that of Trif (2011), whose MATLAB software Chebpack (Trif, 2012) implements a Chebyshev-tau approach allowing the ready solution of a variety linear neutral-type DDEs with discrete and pantograph delay, including higher-order DDEs and DDE eigenvalue problems. The aim of the present work is to describe a unifying framework based on Chebyshev spectral collocation and the barycentric resampling matrix (see Section 2.1) that can solve at once a variety and combination of the types of DDE and FDE problems discussed above, and furthermore in such a way that it can be easily incorporated in Chebfun (Driscoll et al., 2014) to utilise its convenient syntax, nonlinear, and adaptive solves. Rather than attempt to prove convergence results for some limited class of problems, we shall demonstrate by numerous examples the efficacy and generality of the proposed framework.
@@
To that end, the outline of this paper is as follows. In Section 2 we provide an overview of some necessary preliminaries, specifically the barycentric interpolation formula and how it leads to the pseudospectral differentiation and barycentric resampling matrices; our main tools in solving DDEs via spectral collocation. In Section 3 we show how these matrices are combined to solve some simple linear and nonlinear DDEs of various delay types. In Section 4 we discuss what small modifications are required so as to apply the method to FDEs. In Section 5 we provide details of how the approach can be incorporated in the Chebfun/Chebop framework for the automated solution of DDEs/FDEs and demonstrate some more challenging examples solved via this software. In Section 6 we discuss several possible extensions before concluding Section 7.
@@ 2.1 Barycentric interpolation formula
By differentiating the columns of $P(t;\bm{t})$ and evaluating the result at $t=\bm{t}$, one obtains the pseudospectral differentiation matrix, $D=P^{\prime}(\bm{t};\bm{t})$ (Berrut and Trefethen, 2004, Section 9.3). Code for constructing such matrices is given in Table 1. Alternatively, given a different set of points, say $\{\tau_{j}\}_{j=1}^{m}$, $P(\bm{\tau};\bm{t})\bm{y}$ evaluates the degree $n$ polynomial interpolant of the data $\{(t_{k},y_{k})\}_{k=1}^{n}$ at the $m$ points $\bm{\tau}=[\tau_{1},\ldots,\tau_{m}]^{\top}$. Driscoll and Hale (2016) use this barycentric resampling matrix with $m<n$ to downsample discretised linear operators in spectral collocation, leaving space to apply boundary conditions or other side constraints. Here we will use such matrices to evaluate delay terms in DDEs. For instance, a delay term $y(qt-p)$ in a DDE can be discretized as something like $P(q\bm{t}-p;\bm{t})\bm{y}$ [^2] Here we assume $p$ and $q$ are such that $q\bm{t}-p$ is contained within the interpolation interval of (2.1). In general, one may need to subdivide $[0,T]$ and consider multiple interpolants in order to achieve this. See Section 3.2 for details. Concise Matlab code for constructing $P(\bm{\tau};\bm{t})$ is given in Table 1.
@@ 2.2 Interpolation at Chebyshev points
It is well known that interpolation at arbitrary nodes may be ill-conditioned. For example, interpolating at uniformly-spaced nodes leads to the Runge phenomenon and exponential ill-conditioning (Trefethen, 2019, Chapter 13). Suitable nodes for polynomial interpolation cluster towards the endpoints of the interpolation interval. Of such points, the Chebyshev–Gauss–Lobatto nodes,
@@
With these points and weights the barycentric formula (2.1) is numerically stable and the interpolation problem is well-conditioned (Higham, 2004). If the $\{y_{k}\}$ are sampled from a function $y(t)$ analytic in a neighbourhood of $[-1,1]$, then the Chebyshev polynomial interpolant and its derivatives converge geometrically to $y(t)$ as $n$ is increased. In particular, $\|p^{(\nu)}-y^{(\nu)}\|_{L^{\infty}_{[-1,1]}}=\mathcal{O}(\rho^{-n})$ as $n\rightarrow\infty$, where $\rho$ is the radius of the largest Bernstein ellipse with foci at ±$\pm 1$ that lies within the domain of analyticity of $y$ (Trefethen, 2019, Theorem 21.1).
@@ 2.3 Spectral collocation for ODEs
The differentiation matrix, $D$, as computed by Table 1a, is the primary tool for spectral collocation (Trefethen, 2000, Chapter 11). Higher-order differentiation matrices can be computed by taking powers of $D$, although there are more accurate and more efficient approaches of calculating higher-order matrices directly (Baltensperger and Trummer, 2003), (Tee, 2006, (3.28)). Non-constant coefficients can be incorporated by multiplying by a diagonal matrix containing the values of the coefficient function evaluated at the grid points. Nonlinear problems can be solved by a Newton–Raphson iteration, and piecewise smooth problems by domain decomposition (both of which are discussed below). Boundary conditions can be enforced in a number of ways. The most common approach is boundary bordering (sometimes called row replacement), whereby rows of the discretized operator are removed and replaced with the boundary constraints (Boyd, 2001, Section 6.4)(Trefethen, 2000, p.135). An alternative approach is to project the discretized operator onto a lower degree subspace via the barycentric resampling matrix and use the boundary constraints to ‘square-up’ the problem (Driscoll and Hale, 2016). For simplicity we employ the more traditional boundary bordering approach in this work, but any reasonable method of enforcing the boundary conditions (including rectangular projection) is equally applicable.
@@
with solution $y(t)=e^{-t}$. To solve, the $n\times n$ Chebyshev differentiation matrix is constructed using the code in Table 1a and the ODE discretized as $A\bm{y}=\bm{b}$ with $A=D+I$. Boundary conditions are enforced by boundary bordering, i.e., removing the first rows of $A$ and $\bm{b}$ and replacing them with the boundary constraint (here the first row of the identity matrix and the scalar 1, respectively). The resulting square system is then solved for $\bm{y}$, the approximate solution on the Chebyshev grid $\bm{t}$. Concise code for solving (2.6) on the interval $[0,1]$ and a plot demonstrating geometric convergence of the discretisation as $n$ is increased are given in Figure 1.
@@
Refer to captionRefer to caption
Figure 1: Left: MATLAB code for solving the ODE (2.6) using Chebyshev spectral collocation. The line [t,~,w] = chebpts(n, [0,1]) (from Chebfun) returns the n-point Chebyshev grid on the interval $[0,1]$ and the associated barycentric weights (2.4). Centre: Exact and approximate solution using a 12-point discretisation. Right: Geometric convergence of the solution as $n$ is increased down to around machine precision. Here, and throughout (unless otherwise stated), the reported error is the $L^{\infty}$ error at the interpolation points.
@@ 3.1 Proportional delay
We start by considering proportional (sometimes called pantograph-type) delay of the form $y(qt)$, $q<1$ on the interval $[0,1]$. [^3] Any other finite domain can be dealt with via an affine change of variables, and discrete delays introduced by such a change of variables can be dealt with in the manner described in the next section. In this setting $qt\in[0,1]$ and no history function or domain decomposition is required. Such a delay term can therefore be implemented directly by the resampling matrix $P(q\bm{t};\bm{t})$.
@@
with solution $y=e^{-t}$. [^4] A similar DDE appears in (Ali et al., 2009), where it is solved using a Legendre collocation method. A stiff version of the problem is considered by El-Safty and Hussien (1998), where a small parameter is introduced in front of the $u^{\prime}$ term. Discretisation of the differential and identity operators proceeds as in Example 1 and the delay term is discretized using the barycentric resampling matrix as described above. This results now in a system $A\bm{y}=\bm{b}$ where $A=D+I+P(\bm{t}/2;\bm{t})$. As before, boundary conditions can be enforced by row replacement. Code is provided in Figure 2 (left) and, as in the ODE case, we observe geometric convergence as the discretization size is increased (right). Multiple delays of proportional type, e.g., $y(q_{1}t)$, $y(q_{2}t)$, can by included by incorporating additional resampling matrices, $P(q_{1}\bm{t};\bm{t})$, $P(q_{2}\bm{t};\bm{t})$. Extensions to systems of DDEs and/or nonlinear problems proceed as in the ODE case (Driscoll and Hale, 2016, Section 4).
@@
7u = [B;A(2:end,:)]\[y0;b(2:end)];
Refer to captionRefer to caption
Figure 2: Left: MATLAB code for solving the DDE (3.1) using Chebyshev spectral collocation. The code is a simple modification of that from Figure 1 for solving the ODE (2.6). Centre: Solution using 12-point Chebyshev grid. Right: Convergence as $n$ is increased. As in the ODE case, geometric convergence to around the level of machine precision is observed.
@@ 3.2 Discrete delay and domain decomposition
We next consider discrete delay. For simplicity we start only with problems with trivial history functions, i.e., $\phi(t)=0$ for all $t<0$. Non-trivial history functions can be dealt with in the manner described in the next section.
A challenge arising with discrete delays (or with proportional delays on intervals not including the origin) is that discontinuities caused by non-smooth initial data will propagate into the solution whenever the delay becomes active (Baker et al., 1995), and the rapid convergence of the spectral collocation method will be lost (Trefethen, 2000, Chapter 4). If the delay (or delays) are short relative to the length of the solution interval, then many discontinuities will be introduced and the use of a spectral method is unlikely to be beneficial. However, if only a few discontinuities are introduced with locations known a priori, then in these can be dealt with by a domain decomposition/multidomain approach. In particular, if the solution is piecewise analytic on each subdomain, then geometric convergence may be maintained, as in the case of ODEs (Driscoll and Fornberg, 1998). We describe this approach first for a simple problem leading to only one singularity in the domain of interest before outlining the more general approach. A similar approach is employed by (Borgioli et al., 2020) in the context of stability analysis of linear time-periodic delayed dynamical systems (see also Example 11).
@@
From the delay term $y(t-\tfrac{1}{2})$, we know a priori that the solution will contain a jump in its first derivative at $t=\tfrac{1}{2}$ when the delay becomes active (and its second, third, fourth, $\ldots$ derivatives at $t=1,\tfrac{3}{2},2,\ldots$ if we were to integrate further – see below). As such, we discretise on a piecewise domain with two Chebyshev grids, $\bm{t}_{L}\in[0,\tfrac{1}{2}]$ and $\bm{t}_{R}\in[\tfrac{1}{2},1]$, and seek an approximate solution $\bm{y}=[\bm{y}_{L};\bm{y}_{R}]$ on each grid. A discretized differential operator is constructed for each subinterval and a continuity condition $[-\bm{e}_{n_{L}}^{\top},\bm{e}_{1}^{\top}]\bm{y}=0$ enforced so that the solution is continuous at $t=\tfrac{1}{2}$. [^5] For higher-order problems, continuity of additional derivatives is enforced at the interface by including appropriate rows of suitable differentiation matrices. Combined with the initial condition and boundary bordering, this results in a square system that can be solved for $\bm{y}$. Excluding the delay term, details of such domain decomposition are described in detail in (Driscoll and Hale, 2016, Section 4.3).
@@
To incorporate the delay term, we first note that $y(t-\tfrac{1}{2})=0$ for $t<\tfrac{1}{2}$, and so this term may be neglected in the first subinterval. For any $t$ in the second subinterval, $[\tfrac{1}{2},1]$, we note that $t-\tfrac{1}{2}\in[0,\tfrac{1}{2}]$, thus the solution on the first subinterval is interpolated to the appropriate points by the resampling matrix $P(\bm{t}_{R}-\tfrac{1}{2},\bm{t}_{L})$. [^6] If the width and the discretization size of the two intervals is the same, then this will simply be the identity matrix, but in general the resampling matrix will be nontrivial. See Remark 1 for further discussion. This gives rise to the block linear system in the left panel of Figure 3, where again we use boundary bordering to incorpomichiels2020rate the boundary and continuity conditions. For this example, and indeed any initial-value problem DDE, the resulting system is block lower-triangular and one can solve first for $\bm{y}_{L}$ and then for $\bm{y}_{R}$. In the context of solving DDEs, this is referred to as the method of steps. Whilst this structure could certainly be exploited here, FDEs and BVP-DDEs can be implemented using the same framework as described above, and these will not result in block lower-triangular systems. For simplicity and generality, we therefore simply solve the resulting block discretisation all at once using standard direct methods and defer exploration of more efficient domain decomposition strategies for future work. MATLAB code for reproducing the solution to (3.2) (and all other examples in this paper) is available at the online repository: Hale (2024).
$\left[\!\!\!\begin{array}[]{cccc}\bm{e}_{1}^{\top}\\[1.00006pt] \overline{D_{L}}+\overline{I}&\\ -\bm{e}_{n_{L}}^{\top}&\!\!\!\bm{e}_{1}^{\top}\\[1.00006pt] \overline{P}(\bm{t}_{R}-\tfrac{1}{2};\bm{t}_{L})&\!\!\!\overline{D_{R}}+\overline{I}\end{array}\!\!\!\right]\!\left[\!\!\!\begin{array}[]{cccc}\bm{y}_{L}\\ \bm{y}_{R}\end{array}\!\!\!\right]=\left[\!\!\begin{array}[]{cccc}1\\ \bm{0}\\ 0\\ \bm{0}\end{array}\!\!\!\right]$
Refer to captionRefer to caption

### v2

@@ Abstract
A framework for Chebyshev spectral collocation methods for the numerical solution of functional and delay differential equations (FDEs and DDEs) is described. The framework combines interpolation via the barycentric resampling matrix with a multidomain approach used to resolve isolated discontinuities propagated by non-smooth initial data. Geometric convergence in the number of degrees of freedom is demonstrated for several examples of linear and nonlinear FDEs and DDEs with various delay types, including discrete, proportional, continuous, and state-dependent delay. The framework is a natural extension of standard spectral collocation methods and can be readily incorporated into existing spectral discretisations, such as in Chebfun/Chebop, allowing the automated and efficient solution of a wide class of nonlinear functional and delay differential equations.
@@ 1 Introduction
Delay differential equations (DDEs) arise in a variety of application areas, notably population dynamics (Gopalsamy, 1992; Kuang, 1993), epidemiology (Brauer and Castillo-Chávez, 2001; Rihan, 2021), control theory (Kyrychko and Hogan, 2010; Michiels and Niculescu, 2007), and more (Erneux, 2009). Below we review some basic concepts and refer the reader to any of the many excellent introductory texts on DDEs and FDEs for details (Driver, 1977; Hale, 1977; Kolmanovskii and Myshkis, 1999; Michiels and Niculescu, 2007)).
The most elementary DDEs are first-order initial value problems (IVPs) of the form
@@
Like ordinary differential equations (ODEs), most DDEs and FDEs arising in practical settings do not have closed-form solutions, so some kind of approximation is required. Standard numerical methods for the approximation of DDEs like (1.1) typically involve Runge–Kutta-like solvers equipped with a suitable interpolation to evaluate the delay terms (the method of steps), or rewriting the DDE as a system of ODEs with additional variables to represent the delayed terms. For an overview of such methods, we refer the reader to texts such as (Bellen and Zennaro, 2013; Engelborghs et al., 2000b; Shampine and Thompson, 2009), and references therein. A major difficulty arising in DDEs and FDEs not present in ODEs is the propagation of discontinuities arising from non-smooth initial data. Locating and resolving these discontinuities is necessary in order to compute an accurate and efficient approximation, even for low order methods (Baker et al., 1995).
@@
For smooth ODEs, spectral collocation methods (also known as pseudospectral methods) typically exhibit geometric convergence as the number of degrees of freedom is increased, making them particularly efficient when a high degree of accuracy is required (Boyd, 2001; Fornberg, 1996; Trefethen, 2000). Moreover, collocation methods extend naturally to boundary value problems and differential-eigenvalue problems. Given the power of spectral methods in solving ODEs, it is natural to apply them to DDEs and FDEs, and there are numerous examples of this in the literature. We mention a few key papers here, but more detailed overviews of the development of spectral methods in these areas are given by Meng et al. (2019) and Jafari et al. (2021).
@@
One of the first spectral methods for DDEs was introduced by Fox et al. (1971), where a Chebyshev-tau approach was used to solve the linear pantograph equation. Later, a Legendre-tau approach for DDEs with discrete delay was presented by Ito et al. (1991). More recently there have been developments in using spectral methods to solve continuous delay DDEs (Brunner and Hu, 2005, 2007), nonlinear DDEs (Ali et al., 2009), neutral and state-dependent DDEs (Sedaghat et al., 2012), DDE-EVPs (Liu and Ortiz, 1989), and PDDEs (Engelborghs et al., 2000a; Butcher and Bobrenkov, 2011). In addition, spectral collocation methods have proven useful for computing characteristic roots of DDEs (Breda et al., 2005; Wu and Michiels, 2012). Two existing MATLAB software packages closely related to that which we present here are Chebpack (Trif, 2011, 2012) and DDE-BIFTOOL (Engelborghs et al., 2002; Sieber et al., 2016). Chebpack implements a Chebyshev-tau approach allowing the ready solution of a variety of linear neutral-type DDEs with discrete and pantograph delay, including higher-order DDEs and DDE-EVPs. DDE-BIFTOOL, on the other hand, not only solves various forms of DDEs, but is also capable of computing extensive bifurcation analyses for dynamical systems with delay. For PDDEs and DDE-EVPs, DDE-BIFTOOL uses polynomial collocation at Gauss–Legendre nodes (Engelborghs et al., 2000a) in a similar way to that which we describe in Section 3 below.
@@
The present work describes a unifying framework based on Chebyshev spectral collocation and the barycentric resampling matrix (see Section 2.1) that can solve at once a variety and combination of the types of DDE and FDE problems discussed above. The aim is to achieve spectral accuracy using a discretisation that can be easily incorporated in Chebfun (Driscoll et al., 2014) to utilise its convenient syntax and efficient algorithms for nonlinear and adaptive solves. Rather than attempt to rigorously prove convergence results for some limited class of problems, we demonstrate by numerous examples the efficacy and generality of the proposed framework. To that end, the outline of this paper is as follows. In Section 2 we provide an overview of some necessary preliminaries, specifically the barycentric interpolation formula and how it leads to the pseudospectral differentiation and barycentric resampling matrices: our main tools in solving DDEs via spectral collocation. In Section 3 we show how these matrices are combined to solve some simple linear and nonlinear DDEs of various delay types. In Section 4 we discuss what small modifications are required so as to apply the method to FDEs. In Section 5 we provide details of how the approach has been incorporated in the Chebfun/Chebop framework for the automated solution of DDEs/FDEs and demonstrate some more challenging examples solved via this software. In Section 6 we discuss several extensions, including the use of trigonometric interpolants for PDDEs, before concluding in Section 7.
@@ 2.1 Barycentric interpolation formula
By differentiating the columns of $P(t;\bm{t})$ and evaluating the result at $t=\bm{t}$, one obtains the pseudospectral differentiation matrix, $D=P^{\prime}(\bm{t};\bm{t})$ (Berrut and Trefethen, 2004, Section 9.3). Code for constructing such matrices is given in Table 1. Alternatively, given a different set of points, say $\{\tau_{j}\}_{j=1}^{m}$, $P(\bm{\tau};\bm{t})\bm{y}$ evaluates the degree $n$ polynomial interpolant of the data $\{(t_{k},y_{k})\}_{k=1}^{n}$ at the $m$ points $\bm{\tau}=[\tau_{1},\ldots,\tau_{m}]^{\top}$. Driscoll and Hale (2016) use this barycentric resampling matrix with $m<n$ to downsample discretised linear operators in spectral collocation, leaving space to apply boundary conditions or other side constraints. Here we will use such matrices to evaluate delay terms in DDEs. For instance, a delay term $y(qt-p)$ in a DDE can be discretised as something like $P(q\bm{t}-p;\bm{t})\bm{y}$ [^2] Here we assume $p$ and $q$ are such that $q\bm{t}-p$ is contained within the interpolation interval of (2.1). In general, one may need to subdivide $[0,T]$ and consider multiple interpolants in order to achieve this. See Section 3.2 for details. Concise MATLAB code for constructing $P(\bm{\tau};\bm{t})$ is given in Table 1.
@@ 2.2 Interpolation at Chebyshev points
It is well known that interpolation at arbitrary nodes may be ill-conditioned. For example, interpolating at uniformly-spaced nodes leads to the Runge phenomenon and exponential ill-conditioning (Trefethen, 2020, Chapter 13). Suitable nodes for polynomial interpolation cluster towards the endpoints of the interpolation interval. Of such points, the Chebyshev–Gauss–Lobatto nodes,
@@
With these points and weights the barycentric formula (2.1) is numerically stable and the interpolation problem is well-conditioned (Higham, 2004). If the $\{y_{k}\}$ are sampled from a function $y(t)$ analytic in a neighbourhood of $[-1,1]$, then the Chebyshev polynomial interpolant and its derivatives converge geometrically to $y(t)$ as $n$ is increased. In particular, for all $\nu>0$, $\|\rho^{(\nu)}-y^{(\nu)}\|_{L^{\infty}_{[-1,1]}}=\mathcal{O}(\rho^{-n})$ as $n\rightarrow\infty$, where $\rho$ is the radius of the largest Bernstein ellipse with foci at ±$\pm 1$ that lies within the domain of analyticity of $y$ (Trefethen, 2020, Theorem 8.2). Conversely, if $y$ has only $k$ continuous derivatives in $[-1,1]$, then convergence is only algebraic with a rate $\mathcal{O}(n^{-k})$  (Trefethen, 2020, Theorem 7.2).
@@ 2.3 Spectral collocation for ODEs
The differentiation matrix, $D$, as computed by the code in Table 1 (left), is the primary tool for spectral collocation (Trefethen, 2000, Chapter 11). Higher-order differentiation matrices can be computed by taking powers of $D$, although there are more accurate and more efficient approaches of calculating higher-order matrices directly (Baltensperger and Trummer, 2003), (Tee, 2006, (3.28)). Non-constant coefficients can be incorporated by multiplying by a diagonal matrix containing the values of the coefficient function evaluated at the grid points. Nonlinear problems can be solved by a Newton–Raphson iteration, and piecewise smooth problems by domain decomposition (both of which are discussed below). Boundary conditions can be enforced in a number of ways. The most common approach is boundary bordering (sometimes called row replacement), whereby rows of the discretised operator are removed and replaced with the boundary constraints (Boyd, 2001, Section 6.4)(Trefethen, 2000, p.135). An alternative approach is to project the discretised operator onto a lower degree subspace via the barycentric resampling matrix and use the boundary constraints to ‘square-up’ the problem (Driscoll and Hale, 2016). For simplicity we employ the more traditional boundary bordering approach in this work, but any reasonable method of enforcing the boundary conditions (including rectangular projection) is equally applicable.
@@
with solution $y(t)=e^{-t}$. To solve, the $n\times n$ Chebyshev differentiation matrix is constructed using the code in Table 1 (left) and the ODE discretised as $A\bm{y}=\bm{b}$ with $A=D+I$. Boundary conditions are enforced by boundary bordering, i.e., removing the first rows of $A$ and $\bm{b}$ and replacing them with the boundary constraint (here the first row of the identity matrix and the scalar 1, respectively). The resulting square system is then solved for $\bm{y}$, the approximate solution on the Chebyshev grid $\bm{t}$. Concise code for solving (2.6) on the interval $[0,1]$ and a plot demonstrating geometric convergence of the discretisation as $n$ is increased are given in Figure 1.
@@
Figure 1: Left: MATLAB code for solving the ODE (2.6) using Chebyshev spectral collocation. The line [t,~,w] = chebpts(n, [0,1]) (from Chebfun) returns the n-point Chebyshev grid on the interval $[0,1]$ and the associated barycentric weights (2.4). Centre: Exact solution (solid) and approximate solution (dots) using an $n=$ 12-point discretisation. Right: The solid line shows geometric convergence of the solution down to around machine precision, $M_{\varepsilon}=2^{-52}$, as $n$ is increased. Here, and throughout, the reported error is the $\ell^{\infty}$ error at the interpolation points. The dashed line shows (again, here and throughout) $M_{\varepsilon}\times{\rm cond}(A)$ as a function of $n$.
@@ 3.1 Proportional delay
We start by considering proportional (sometimes called pantograph-type) delay of the form $y(qt)$, $q<1$ on the interval $[0,1]$. [^3] Any other finite domain can be dealt with via an affine change of variables as in (2.5). Discrete delays introduced by such a change of variables can be dealt with in the manner described in the next section. In this setting $qt\in[0,1]$ and no history function or domain decomposition is required. Such a delay term can therefore be implemented directly by the resampling matrix $P(q\bm{t};\bm{t})$.
@@
with solution $y=e^{-t}$. [^4] A similar DDE appears in (Ali et al., 2009), where it is solved using a Legendre collocation method. A stiff version of the problem is considered by El-Safty and Hussien (1998), where a small parameter is introduced in front of the $y^{\prime}$ term. Discretisation of the differential and identity operators proceeds as in Example 1 and the delay term is discretised using the barycentric resampling matrix as described above. This results now in a system $A\bm{y}=\bm{b}$ where $A=D+I+P(\bm{t}/2;\bm{t})$. As before, boundary conditions can be enforced by row replacement. Code is provided in Figure 2 (left) and, as in the ODE case, we observe geometric convergence as the discretisation size is increased (right). Multiple delays of proportional type, e.g., $y(q_{1}t)$, $y(q_{2}t)$, can by included by incorporating additional resampling matrices, $P(q_{1}\bm{t};\bm{t})$, $P(q_{2}\bm{t};\bm{t})$. Extensions to systems of DDEs and/or nonlinear problems proceed as in the ODE case—see (Driscoll and Hale, 2016, Section 4).
@@
7y = [B;A(2:end,:)]\[y0;b(2:end)];
Figure 2: Left: MATLAB code for solving the DDE (3.1) using Chebyshev spectral collocation. The code is a simple modification of that from Figure 1 for solving the ODE (2.6). Centre: Solution using a 12-point Chebyshev grid. Right: Convergence of the solution (solid) and $M_{\varepsilon}\times{\rm cond}(A)$ (dashed) as $n$ is increased. As in the ODE case, geometric convergence to around the level of machine precision is observed. The condition number of the discretised DDE behaves similarly to the ODE case.
@@ 3.2 Discrete delay and domain decomposition
We next consider discrete delay. For simplicity we start with problems that have only trivial history functions, i.e., $\phi(t)=0$ for all $t<0$. Non-trivial history functions can be dealt with in the manner described in the next section.
A challenge arising with discrete delays (or with proportional delays on intervals not including the origin) is that discontinuities caused by non-smooth initial data will propagate into the solution whenever the delay becomes active (Baker et al., 1995), and the rapid convergence of the spectral collocation method will be lost (Trefethen, 2000, Chapter 4). If the delay (or delays) are short relative to the length of the solution interval, then many discontinuities will be introduced and the use of a spectral method is unlikely to be beneficial. However, if only a few discontinuities are introduced with locations known a priori, then these can be dealt with by a domain decomposition/multidomain approach. In particular, if the solution is piecewise analytic on each subdomain, then geometric convergence may be maintained, as in the case of ODEs (Driscoll and Fornberg, 1998). We describe this approach first for a simple problem leading to only one singularity in the domain of interest before outlining the more general approach. A similar approach is employed by (Borgioli et al., 2020) and (Engelborghs et al., 2000a) in the context of stability analysis of time-periodic delayed dynamical systems (see also Example 11).
@@
From the delay term $y(t-\tfrac{1}{2})$, we know a priori that the solution will contain a jump in its first derivative at $t=\tfrac{1}{2}$ when the delay becomes active (and its second, third, fourth, $\ldots$ derivatives at $t=1,\tfrac{3}{2},2,\ldots$ if we were to integrate further—see below). As such, we discretise on a piecewise domain with two Chebyshev grids, an $n_{L}$-point grid, $\bm{t}_{L}$, on $[0,\tfrac{1}{2}]$ and an $n_{R}$-point grid, $\bm{t}_{R}$, on $[\tfrac{1}{2},1]$, and seek an approximate solution $\bm{y}=[\bm{y}_{L};\bm{y}_{R}]$ on each grid. A discretised differential operator is constructed for each subinterval and a continuity condition $[-\bm{e}_{n_{L},n_{L}}^{\top},\bm{e}_{n_{R},1}^{\top}]\bm{y}=0$ (where $\bm{e}_{n,k}$ denotes the $k$th column of an $n\times n$ identity matrix) enforced so that the solution is continuous at $t=\tfrac{1}{2}$. [^5] For higher-order problems, continuity of additional derivatives is enforced at the interface by including appropriate rows of suitable differentiation matrices. Combined with the initial condition and boundary bordering, this results in a square system that can be solved for $\bm{y}$. Excluding the delay term, details of such domain decomposition are described in detail in (Driscoll and Hale, 2016, Section 4.3).
@@
To incorporate the delay term, we first note that $y(t-\tfrac{1}{2})=0$ for $t<\tfrac{1}{2}$, so this term may be neglected in the first subinterval. For any $t$ in the second subinterval, $[\tfrac{1}{2},1]$, we note that $t-\tfrac{1}{2}\in[0,\tfrac{1}{2}]$; thus the solution on the first subinterval is interpolated to the appropriate points by the resampling matrix $P(\bm{t}_{R}-\tfrac{1}{2},\bm{t}_{L})$. [^6] If the width and discretisation size of the two intervals is the same, then this will simply be the identity matrix, but in general the resampling matrix will be nontrivial. See Remark 1 for further discussion. This gives rise to the block linear system in the left panel of Figure 3, where again we use boundary bordering to incorporate the boundary and continuity conditions. For this example, and indeed any IVP-DDE, the resulting system is block lower triangular (BLT) and one can solve first for $\bm{y}_{L}$ and then for $\bm{y}_{R}$. In the context of solving DDEs, this is referred to as the method of steps (Hale, 1977, Chapter 1.2). Whilst this structure could certainly be exploited here, FDEs and boundary-value problem DDES (BVP-DDEs) can be implemented using the same framework as described above, and these will not result in block lower triangular systems. Therefore, for simplicity and generality, we solve the whole resulting block discretisation at once using standard direct methods and defer exploration of more efficient domain decomposition strategies for future work. MATLAB code for reproducing the solution to (3.2) (and all other examples in this paper) is available at the online repository (Hale, 2024).
$\left[\!\!\!\begin{array}[]{cccc}\bm{e}_{n_{L},1}^{\top}\\[1.00006pt] \overline{D_{L}}+\overline{I}&\\ -\bm{e}_{n_{L},n_{L}}^{\top}&\!\!\!\bm{e}_{n_{R},1}^{\top}\\[1.00006pt] \overline{P}(\bm{t}_{R}-\tfrac{1}{2};\bm{t}_{L})&\!\!\!\overline{D_{R}}+\overline{I}\end{array}\!\!\!\right]\!\left[\!\!\!\begin{array}[]{cccc}\bm{y}_{L}\\ \bm{y}_{R}\end{array}\!\!\!\right]=\left[\!\!\begin{array}[]{cccc}1\\ \bm{0}\\ 0\\ \bm{0}\end{array}\!\!\!\right]$